# Tunable transport in bi-disperse porous materials with vascular structure

### v1

@@ Abstract
We study transport in synthetic, bi-disperse porous structures, with arrays of microchannels interconnected by a nanoporous layer. These structures are inspired by the xylem tissue in vascular plants, in which sap water travels from the roots to the leaves to maintain hydration and carry micronutrients. We experimentally evaluate transport in three conditions: high pressure-driven flow, spontaneous imbibition, and transpiration-driven flow. The latter case resembles the situation in a living plant, where bulk liquid water is transported upwards in a metastable state (negative pressure), driven by evaporation in the leaves; here we report stable, transpiration-driven flows down to $\sim-15$ MPa of driving force. By varying the shape of the microchannels, we show that we can tune the rate of these transport processes in a predictable manner, using a simple analytical approach and numerical simulations of the flow field in the bi-disperse media. We also show that the spontaneous imbibition behavior of a single structure – with fixed geometry – can behave very differently depending on its preparation (filled with air, vs. evacuated), because of a dramatic change in the conductance of vapor in the microchannels; this change offers a second way to tune the rate of transport in bi-disperse, xylem-like structures, by switching between air-filled and evacuated states.
@@ II.1 Samples
We fabricated xylem-like, composite structures with regularly spaced microchannels interconnected through a nanoporous matrix by assembling (through anodic bonding) a nanoporous silicon layer (poSi) etched into the surface of a silicon wafer by anodization to glass patterned by photolithography (Figure 2). Details on the fabrication methods can be found elsewhere [14, 17].
@@
Figure 2a presents the geometrical properties of the xylem-like designs. Figure 2b presents micrographs of two of the geometries studied. Table 1 provides parameter values of all designs. In all designs, the poSi layer had a thickness, $H_{\mathrm{p}}=5\,\mathrm{\mu m}$ and contained interconnected pores of radius, $r_{\mathrm{p}}\simeq 1.5$ nm (as characterized previously in Vincent et al. [14, 17, 21]). The microchannel segments had a depth, $H_{\mathrm{v}}=32\,\mathrm{\mu m}$ (except for the Vein sample, $H_{\mathrm{v}}\simeq 20\,\mathrm{\mu m}$) and various in-plane lengths ($\ell$) and widths ($w$). These channels were arranged in various regular, rectilinear arrays with the same edge-to-edge separation ($2d$) in both principal directions ($x$ and $y$); the value of $d$ varied with the lateral dimensions to maintain a similar areal fraction of the vessels across designs: $\gamma\sim 0.3$. A sample with no vessels (Blank) was also prepared.
@@ II.2 Experiments
Figure 3: Different types of experiments performed on an individual sample. (a) High pressure flow, with measurement of flow rate using the position, $X(t)$, of the meniscus in the serpentine channel or in the distribution channel (horizontal). (b) Transpiration-driven emptying in a sub-saturated vapor, with a similar measurement of the flow rate using $X(t)$. (c) Spontaneous imbibition triggered by capillary condensation in the nanoporous matrix, with the microstructure either filled with air (diffusive transport of water vapor), or evacuated (convective transport of water vapor); $Y(t)$ corresponds to the position of the wetting front, which we measure optically. Large, orange arrows indicate the direction of the flow, dark brown shading indicates wet nanoporous matrix, and light brown shading indicates dry nanoporous matrix.
@@
We performed two types of imbibition experiments. For the first type, we first dried the samples for at least 48 hours in dry air (humidity $6$% RH). Then, we placed the samples in a closed box with a humidity of $93$% RH, set by an unsaturated solution of sodium chloride (NaCl). For the second type, the samples were evacuated for at least 24 hours in a vacuum chamber. Then, we released water vapor in the chamber to impose a fixed vapor pressure of $p=0.93\,p_{\mathrm{sat}}$ ($93$% RH). As a result, imbibition occurred with air-filled nanopores and channels in the first type of experiments and with evacuated nanopores and channels in the second type.
Note that during these imbibition experiments, the microchannels never filled with liquid water, because bulk liquid water was not the stable phase for $p<p_{\mathrm{sat}}$. Thus, the microstructure only filled with water vapor. On the other hand, liquid phase can be stable under these conditions in confinement [23, 24], so that imbibition could proceed in the liquid phase in the nanoporous layer [21].
@@ III Theory
III.1 Basic equations
Transport
@@ III.1 Basic equations
Transport mechanisms in the microchannels vary depending on the situation. In the cases in Figure 3a-b (high pressure and transpiration), the microstructure was filled with liquid and transport occurred through pressure-driven, Poiseuille-like flow. Since the depth, $H_{\mathrm{v}}$ of the microchannels was small compared to their other dimensions, we evaluate transport dynamics using the expression of planar Poiseuille flow for the mass flux density (mass flow per unit cross-sectional area, $q_{\ell}\,\mathrm{[kg.m^{-2}.s^{-1}]}$):
@@
In imbibition experiments, however, transport in the microchannels occurred in the vapor phase, either by diffusion through air (air-filled state, Figure 3c) or as a pressure-driven (convective) flow of pure water vapor (evacuated state, Figure 3c). From Fick’s law, transport by diffusion leads to the following prediction of the mass flux density:
@@
Kelvin equation
@@
$P=P_{\mathrm{ref}}+\frac{RT}{v_{\mathrm{m}}}\ln\left(\frac{p}{p_{\mathrm{sat}}}\right)=P_{\mathrm{ref}}-\Psi(p)$ (6)
@@
III.2 Relative magnitudes of transport mechanisms
@@
For high pressure flow (Figure 3a) and transpiration (Figure 3b) for which the microchannels were filled with water, we predict (Equations (2) and (3)) the ratio of fluxes in the channels ($q_{\ell}$) and in the nanoporous layer ($q_{\mathrm{po}}$) to be:
@@
For imbibition with evacuated microchannels (Figure 3c), we predict the ratio of vapor flux in the channels ($q_{\mathrm{v,vac}}$) to the Darcy flux in the poSi ($q_{\mathrm{po}}$) to be:
@@
Finally, for imbibition with air-filled microchannels (Figure 3c), we use Equation (4) and the derivative of Equation (6) to predict the flux in air-filled microchannels ($q_{\mathrm{v,air}}$) and find the ratio:
@@
III.3 Effective permeability of the composite structures
@@
In our numerical solution, we divided the unit cell into two domains: the microchannel (inside white dotted line in Figure 4a) and the nanoporous domain (outside white dotted line). Consistent with its high conductance (Equations (9) and (10)), we treated the microchannel as a "lumped capacitance" with a single, uniform pressure at each time step. In the porous domain, we solved the poroelastic diffusion equation using an explicit finite difference scheme. The boundary conditions imposed on the porous domain were no flux (right and left), fixed pressure $P_{\mathrm{in}}$ at the bottom, fixed pressure $P_{\mathrm{out}}$ on top, and uniform pressure along the boundary between the microchannel and the rest of the domain (Figure 4a, white dotted line). We have provided a detailed description of this numerical approach elsewhere (see Supplementary Information of [14], and section 4.2.1 of [32]). We allowed the solution to reach steady state for a fixed difference in pressure imposed on the left ($P_{\mathrm{in}}$) and right boundaries ($P_{\mathrm{out}}$) to assess the steady-state mass flow rate, $Q$ and thus the effective permeability increase from Equation (15).
We numerically measured $\kappa/\kappa_{0}$ for a variety of geometries with different surface coverage ($\gamma$) of the microstructure and with varying aspect ratio of the microchannels (see Figure 4b, symbols). The $\gamma=0.3$ case is close to our experimental samples (see Table 1); however our fabrication procedure induced variations in the values of $\gamma$ for the different samples. Consequently, we also ran specific numerical simulations with each actual sample geometry, as measured on optical micrographs; the corresponding values are indicated in the last column in Table 1.
@@
III.4 Expressions corresponding to the different experimental situations
@@
High pressure experiments (Figure 3a) correspond to a steady-state situation with an external pressure, $P$, pushing the liquid towards the serpentine channel (at pressure, $P_{\mathrm{s}}$). The typical magnitude of $P$ was $\sim 10$ MPa ($10^{7}$ Pa). Since the sample was evacuated before high pressure experiments, the gas pressure in the bubble at the end of the serpentine channel was the saturation pressure of water: $P_{\mathrm{s}}=p_{\mathrm{sat}}(T)\simeq 2$ kPa $\sim 10^{3}$ Pa (we can neglect here the effect of curvature on the vapor pressure since the meniscus dimensions are in the micrometer scale). Also, the Laplace pressure due to the curvature of the liquid-vapor interface between bubble and liquid in the channel is on the order of $2\sigma/H_{\mathrm{p}}\simeq 4$ kPa. Thus, $P_{\mathrm{s}}\sim 10^{3}\,\mathrm{Pa}\ll P\sim 10^{7}\,\mathrm{Pa}$ so that the pressure difference between the edge of the sample and the serpentine channel is, to very good approximation, $\Delta P=-P$. From Darcy’s law (Equation 14), it follows that the mass flux is
@@
In the situation of Figure 3c, water vapor condenses at the open edge and the condensed liquid in the nanopores invades the structure, driven by the capillary pressure, $\Delta P_{\mathrm{c}}=2\sigma\cos\theta/r_{\mathrm{p}}$, where $r_{\mathrm{p}}$ is the nanopores radius and $\theta$ the contact angle of water on the pore walls. This situation is similar to regular spontaneous imbibition, i.e., capillary invasion when the sample edge is soaked in bulk liquid water; this process is described by the Lucas-Washburn equation [35, 36]. However, in the present situation where imbibition is triggered by capillary condensation, the liquid-vapor equilibrium at the open edge introduces another capillary pressure, $\Psi(p)$ (through the Kelvin equation, see Equation 6) that resists the invasion [21]. As a result the invasion dynamics (front position, $Y(t)$) is described by a modified Lucas-Washburn equation that takes into account the competition of these two capillary pressures, as we have shown previously [21]:
@@
where $\phi$ is the porosity of the nanoporous layer; $\omega=2\kappa(\Psi-\Delta P_{\mathrm{c}})/\phi=Y^{2}/t$ is the Lucas-Washburn "velocity" of imbibition (homogeneous to a diffusivity $\mathrm{[m^{2}/s]}$) [21], and depends on the humidity imposed around the sample through $\Psi(p)$.

### v2

@@ Abstract
We study transport in synthetic, bi-disperse porous structures, with arrays of microchannels interconnected by a nanoporous layer. These structures are inspired by the xylem tissue in vascular plants, in which sap water travels from the roots to the leaves to maintain hydration and carry micronutrients. We experimentally evaluate transport in three conditions: high pressure-driven flow, spontaneous imbibition, and transpiration-driven flow. The latter case resembles the situation in a living plant, where bulk liquid water is transported upwards in a metastable state (negative pressure), driven by evaporation in the leaves; here we report stable, transpiration-driven flows down to $\sim-15$ MPa of driving force. By varying the shape of the microchannels, we show that we can tune the rate of these transport processes in a predictable manner, using a simple analytical (effective medium) approach and numerical simulations of the flow field in the bi-disperse media. We also show that the spontaneous imbibition behavior of a single structure – with fixed geometry – can behave very differently depending on its preparation (filled with air, vs. evacuated), because of a dramatic change in the conductance of vapor in the microchannels; this change offers a second way to tune the rate of transport in bi-disperse, xylem-like structures, by switching between air-filled and evacuated states.
@@ II.1 Samples
We fabricated xylem-like, composite structures with regularly spaced microchannels interconnected through a nanoporous matrix by assembling (through anodic bonding) two surfaces. The first surface consisted of a nanoporous silicon layer (poSi) etched into the surface of a silicon wafer by anodization. The second surface was glass, patterned with microchannels by photolithography (Figure 2). Details on the fabrication methods can be found elsewhere [14, 17]. Separating the fabrication of the nanoporous layer and the microchannels in two separate surfaces results in simpler fabrication procedures compared to creating all structures into the silicon wafer (e.g. wet etching instead of reactive ion etching, parallel instead of sequential processing etc.), but the latter is also possible [16]. We do not expect these two fabrication procedures to result in significantly different behavior, as the final layers remain essentially two-dimensional (nanoporous layer and microchannel depths small compared to other system dimensions).
@@
Figure 2a presents the geometrical properties of the xylem-like designs. Figure 2b presents micrographs of two of the geometries studied. Table 1 provides parameter values of all designs. In all designs, the poSi layer had a thickness, $H_{\mathrm{p}}=5\,\mathrm{\mu m}$ and contained interconnected pores of radius, $r_{\mathrm{p}}\simeq 1.5$ nm (as characterized previously in Vincent et al. [14, 17, 21]). The microchannel segments had a depth, $H_{\mathrm{v}}=32\,\mathrm{\mu m}$ (except for the Vein sample, $H_{\mathrm{v}}\simeq 20\,\mathrm{\mu m}$) and various in-plane lengths ($\ell$) and widths ($w$). These channels were arranged in various regular, rectilinear arrays with the same edge-to-edge separation ($2d$) in both principal directions ($x$ and $y$); the value of $d$ varied with the lateral dimensions to maintain a similar areal fraction of the microchannels across designs: $\gamma\sim 0.3$. A sample with no microchannels (Blank) was also prepared.
@@ II.2 Experiments
Figure 3: Different types of experiments performed on an individual sample. (a) High pressure flow, with measurement of flow rate using the position, $X(t)$, of the meniscus in the serpentine channel or in the distribution channel (horizontal). (b) Transpiration-driven emptying in a sub-saturated vapor, with a similar measurement of the flow rate using $X(t)$. (c) Spontaneous imbibition triggered by capillary condensation in the nanoporous matrix, with the microstructure either filled with air (diffusive transport of water vapor, represented by the blue, broken arrow), or evacuated (convective transport of water vapor, parallel blue arrows, corresponding to a Poiseuille flow developed in the $z$ direction, see Figure 4c); $Y(t)$ corresponds to the position of the wetting front, which we measure optically. Large, orange arrows indicate the direction of the flow, dark brown shading indicates wet nanoporous matrix, and light brown shading indicates dry nanoporous matrix.
@@
We performed two types of imbibition experiments. For the first type, we first dried the samples for at least 48 hours in dry air (humidity $6$% RH). Then, we placed the samples in a closed box with a humidity of $93$% RH, set by an unsaturated solution of sodium chloride (NaCl). For the second type, the samples were evacuated for at least 24 hours in a vacuum chamber. Then, we released water vapor in the chamber to impose a fixed vapor pressure of $p=0.93\,p_{\mathrm{sat}}$ ($93$% RH). As a result, imbibition occurred with air-filled nanopores and microchannels in the first type of experiments and with evacuated nanopores and microchannels in the second type.
Note that during all imbibition experiments, the microchannels never filled with liquid water, because the humidity used to trigger condensation in the nanopores was not sufficiently high to induce condensation in the microchannels, nor capillary suction from the nanopores into the microchannel (see Theory, section III.2)
@@ III Theory
III.1 Hypotheses
Since dimensions of the structures in the $z$-direction ($H_{\mathrm{p}}$ for the nanoporous layer, $H_{\mathrm{v}}$ for the microchannels) are much smaller than in other directions, we assume that all gradients leading to transport in the system are in the $x$ and $y$ directions (Figure 2a), and also that there is fast exchange between the microchannels and nanoporous layer in the $z$ direction (local equilibrium).
III.2 Basic equations
Transport in the nanopores.
@@ III.2 Basic equations
Transport in the microchannels.
Figure 4: Sketches of transport mechanisms along the length ($y$) of the microchannels in three situations: (a) channel filled with liquid (Poiseuille flow of liquid water), (b) channel filled with air (diffusive transport of water vapor), and (c) evacuated channel (Poiseuille flow of water vapor). The sketches depict a cross-section along the length of a microchannel. The poSi layer is represented as interconnected pores, drawn with exagerated diameter compared to their actual size. The squiggle arrows in (b-c) represent evaporation and condensation at the left and right of the channel, respectively. The situation in (a) corresponds to high-pressure or transpiration experiments (Figure 3a-b). The situation in (b) and (c) correspond to spontaneous imbibition experiments (Figure 3c), in the air-filled and evacuated cases respectively. In all cases, transport in the microchannel occurs in parallel with a Darcy flow in the poSi layer, driven by a pressure gradient (hence, the increasing meniscus curvature from left to right in panels b-c), see Equation (2). Situations described in panels (a), (b) and (c) correspond to Equations (3), (4) and (5), respectively.
In regions of the sample where microchannels are in contact with the porous layer, transport in the microchannels can occur in parallel to Darcy flow in the nanopores (Figure 4). This process happens either by division of the liquid flow when the microchannels are filled with liquid (Figure 4a), or by evaporation, transport through the vapor phase, and condensation when the channels are air-filled or evacuated (Figure 4a).
In the cases in Figure 3a-b (high pressure and transpiration), the microstructure was filled with liquid and transport occurred through pressure-driven, Poiseuille-like flow (Figure 4a). Since the depth, $H_{\mathrm{v}}$ of the microchannels was small compared to their other dimensions, we evaluate transport dynamics using the expression of planar Poiseuille flow for the mass flux density (mass flow per unit cross-sectional area, $q_{\ell}\,\mathrm{[kg.m^{-2}.s^{-1}]}$):
@@
In imbibition experiments (Figure 3c), however, transport in the microchannels occurred in the vapor phase, either by diffusion through air (air-filled state, Figure 4b) or as a pressure-driven (convective) flow of pure water vapor (evacuated state, Figure 4c). From Fick’s law, transport by diffusion leads to the following prediction of the mass flux density:
@@
Liquid-vapor equilibrium (Kelvin equation).
@@
$P=P_{\mathrm{ref}}+\frac{RT}{v_{\mathrm{m}}}\ln\left(\frac{p}{p_{\mathrm{sat}}}\right)=P_{\mathrm{ref}}+\Psi(p)$ (6)
@@
Capillary condensation and spontaneous filling.
Kelvin equation (6) can also be used to estimate the humidity at which pores or channels fill spontaneously by capillary condensation, or by meniscus aspiration from the porous medium. In order to fill the hydrophilic microchannel, one must form a meniscus in the channel, resulting in a reduced pressure in the liquid due to capillarity ($P\sim-2\sigma/H_{\mathrm{v}}\simeq-5$ kPa from Laplace’s law, with $\sigma=0.073$ N/m, the surface tension of the liquid). Condensation becomes possible when this capillary pressure is compatible with Kelvin’s equation (6), in other words the typical relative humidity necessary for capillary condensation in the microchannels is $p/p_{\mathrm{sat}}=\exp\left(-2\sigma v_{\mathrm{m}}/(RTH_{\mathrm{v}})\right)\simeq 0.99997$, neglecting $P_{\mathrm{ref}}$. In other words, condensation in the microchannels is possible only if the imposed relative humidity is extremely close to $100$% RH, which was not the case in our experiments. Similarly, Kelvin’s equation implies that the pressure in the liquid in the nanopores during imbibition experiments varies between $\Psi(p)\simeq-10$ MPa at the edge of the sample, down to $\Delta P_{\mathrm{c}}\simeq-70$ MPa at the imbibition front (see below, section III.5). Since these values are much larger in magnitude than the microchannel capillary pressure ($\simeq-5$ kPa), spontaneous invasion of the microchannels was impossible in our experiments, due to the much stronger suction of the nanoporous layer.
III.3 Relative magnitudes of transport mechanisms
@@
Because the cross-sections available for flow are similar in the poSi layer and in the microchannels ($H_{\mathrm{p}}$ comparable to $H_{\mathrm{v}}$), we assume below that the ratio of mass flux densities in the porous layer and in the microchannel is a reasonable measure of relative transport between these two elements.
For high pressure flow (Figure 3a) and transpiration (Figure 3b) for which the microchannels were filled with water (Figure 4a), we predict (Equations (2) and (3)) the ratio of fluxes in the channels ($q_{\ell}$) and in the nanoporous layer ($q_{\mathrm{po}}$) to be:
@@
For imbibition with evacuated microchannels (Figure 3c, Figure 4c), we predict the ratio of vapor flux in the channels ($q_{\mathrm{v,vac}}$) to the Darcy flux in the poSi ($q_{\mathrm{po}}$) to be:
@@
Finally, for imbibition with air-filled microchannels (Figure 3c, Figure 4b), we use Equation (4) and the derivative of Equation (6) to predict the flux in air-filled microchannels ($q_{\mathrm{v,air}}$) and find the ratio:
@@
Since transport in the microchannels is either very large (Equations 9-10) or very small (Equation 11) compared to that in the nanoporous layer, a detailed description of the local exchange and distribution of flow between microchannels and the nanopores situated below in the z-direction (Figure 4) is not necessary: in the following, we will simply assume that microchannels either completely bypass the nanoporous layer below it when filled with liquid or water vapor, and are invisible (all flow in the nanoporous layer below) when filled with air.
@@
Note that varying the ratio $H_{\mathrm{v}}/H_{\mathrm{p}}$ is in principle another way to tune the relative flow between microchannels and porous layer. For example, from Equation (10), the mass flux density decreases rapidly in the evacuated microchannel when its height is decreased, thus by lowering this value enough, one should be able to make it similar to the mass flux density in the porous layer; this choice would result in a similar behavior between evacuated and air-filled samples (negligible flow in the microchannels). We have not explored this possibility in the present study. This observation shows the importance of the choice of physical and geometrical parameters for both nanoporous layer and microchannel for the switching and tunable behavior to be possible.
III.4 Effective permeability of the composite structures
@@
In our numerical solution, we divided the unit cell into two domains: the microchannel (inside white dotted line in Figure 5a) and the nanoporous domain (outside white dotted line). Consistent with its high conductance (Equations (9) and (10)), we treated the microchannel as a "lumped capacitance" with a single, uniform pressure at each time step. In the porous domain, we solved the poroelastic diffusion equation using an explicit finite difference scheme. The boundary conditions imposed on the porous domain were no flux (right and left), fixed pressure $P_{\mathrm{in}}$ at the bottom, fixed pressure $P_{\mathrm{out}}$ on top, and uniform pressure along the boundary between the microchannel and the rest of the domain (Figure 5a, white dotted line). We have provided a detailed description of this numerical approach elsewhere (see Supplementary Information of [14], and section 4.2.1 of [30]). We allowed the solution to reach steady state for a fixed difference in pressure imposed on the bottom ($P_{\mathrm{in}}$) and top boundaries ($P_{\mathrm{out}}$) to assess the steady-state mass flow rate, $Q$ and thus the effective permeability increase from Equation (15).
We numerically calculated $\kappa/\kappa_{0}$ for a variety of geometries with different surface coverage ($\gamma$) of the microstructure and with varying aspect ratio of the microchannels (see Figure 5b, symbols). The $\gamma=0.3$ case is close to our experimental samples (see Table 1); however our fabrication procedure induced variations in the values of $\gamma$ for the different samples. Consequently, we also ran specific numerical simulations with each actual sample geometry, as measured on optical micrographs; the corresponding values are indicated in the last column in Table 1.
@@
III.5 Expressions corresponding to the different experimental situations
@@
High pressure experiments (Figure 3a) correspond to a steady-state situation with an external pressure, $P$, pushing the liquid towards the serpentine channel (at pressure, $P_{\mathrm{s}}$). The typical magnitude of $P$ was $\sim 10$ MPa ($10^{7}$ Pa). Since the sample was evacuated before high pressure experiments, the gas pressure in the bubble at the end of the serpentine channel was the saturation pressure of water: $P_{\mathrm{s}}=p_{\mathrm{sat}}(T)\simeq 2$ kPa $\sim 10^{3}$ Pa (we can neglect here the effect of curvature on the vapor pressure since the meniscus dimensions are in the micrometer scale). Also, the Laplace pressure due to the curvature of the liquid-vapor interface between bubble and liquid in the channel is on the order of $-2\sigma/H_{\mathrm{v}}\simeq-5$ kPa. Thus, $P_{\mathrm{s}}\sim 10^{3}\,\mathrm{Pa}\ll P\sim 10^{7}\,\mathrm{Pa}$ so that the pressure difference between the edge of the sample and the serpentine channel is, to very good approximation, $\Delta P=-P$. From Darcy’s law (Equation 14), it follows that the mass flux is
@@
In the situation of Figure 3c, water vapor condenses at the open edge and the condensed liquid in the nanopores invades the structure, driven by the capillary pressure, $\Delta P_{\mathrm{c}}=-2\sigma\cos\theta/r_{\mathrm{p}}$, where $r_{\mathrm{p}}$ is the nanopores radius and $\theta$ the contact angle of water on the pore walls. From our previous studies [17, 21], $\Delta P_{\mathrm{c}}\simeq-70$ MPa for the nanoporous silicon layer. This situation is similar to regular spontaneous imbibition, i.e., capillary invasion when the sample edge is soaked in bulk liquid water; this process is described by the Lucas-Washburn equation [33, 34]. However, in the present situation where imbibition is triggered by capillary condensation, the liquid-vapor equilibrium at the open edge introduces another capillary pressure, $\Psi(p)$ (through the Kelvin equation, see Equation 6) that resists the invasion [21]. As a result the invasion dynamics (front position, $Y(t)$) is described by a modified Lucas-Washburn equation that takes into account the competition of these two capillary pressures, as we have shown previously [21]:
@@
where $\phi$ is the porosity of the nanoporous layer; $\omega=2\kappa(\Psi-\Delta P_{\mathrm{c}})/\phi=Y^{2}/t$ is the Lucas-Washburn "velocity" of imbibition (homogeneous to a diffusivity $\mathrm{[m^{2}/s]}$) [21], and depends on the humidity imposed around the sample through $\Psi(p)$. In our experiments, we used a fixed value of relative humidity, $p/p_{\mathrm{sat}}=0.93$ (see Methods, section II.2), resulting in $\Psi(p)\simeq-9.9$ MPa from Equation (6). As a result, the imbibition flow was driven by a gradient of pressure going from $\Psi(p)\simeq-9.9$ MPa at the sample edge, down to $\Delta P_{\mathrm{c}}\simeq-70$ MPa at the imbibition front.